CALCULATE AVERAGE SPEED IN KM/H OF CARLOS! WHO RUNS TO THE STORE WHICH IS 4 KM AWAY IN 30 MINUTES?
Answer by Gogonati(855)

30 min= 1/2 h. The average speed of Carlos is v=d/t=4 km/(1/2)h=8 km/h.